Prev Next

SW&IT Meeting Minutes 2019-06-03

Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

20190520
LR20190520 - Bec, Matthieu 68.0
LR20190520 - Colson, Andrew 80.0
LR20190520 - Edmeades, David 64.0
LR20190520 - Sola, Igor 40.0
LR20190520 - Kubanek, Petr 92.0
LR20190520 - Hooper, Stephen 64.0
LR20190520 - Zhang, Yang 64.0
LR20190520 - Zhang, Xianyu 69.0

News from management

  • yearly evaluations

Previous Action Items

  • announcing and delivering of new s/w

Open topics

Team reports

MB

  • OC, WPC, Management
  • dms x/y, psd plot
  • review vibration requirements for PG, contacted vendors for remote sensors
  • obs1/fedora30 activities, packaging, etc.
  • Ketiv / Bucki telecon
next
  • WPC obs1 update
  • PE2018

DE

  • Final edits to L4AO site and registration spreadsheet
  • LBTO site redesign mockup
    • Started a document to note issues that I've noticed in this first run-through
  • Fresh UITS colocation quote obtained
    • We may have an internal switch as long as it doesn't connect to UITS gear
    • We pay for data connections to PDUs
    • They estimate ~4 hours for rack and stack, pending a rack diagram
  • STSCI DSS copy to Synology, ~25h
  • Ordered USB sticks for our stock
  • New 4D server on network
    • The jumper Ethernet cable between the two servers and the associated network config does appear to be critical to the software. Reverting it to the as-shipped config stopped the software crashes Mike reported.
  • Rebuilding failed 8TB disk on passive Synology node
Next:
  • Matlab licensing Tucson/Mountain
  • 4D mount requested by Xianyu

PK

  • Ice3.7 problems - IIF is probably fixed, AOS awaits; need to test on mountain hopefully next week
  • corrupted shared memory tentatively related to gshmserver running on other TCS machines
  • start/stop redesign - to allow for TCS release switch
  • AGW SSD planning
next:
  • force shutdown of gshmserver (to make sure we will be able to switch releases without need to restart)
  • finish Ice3.7 changes

XZ

Last week
  • OVMSTT translationmotion draft review and dicussion
  • Implement 4D interface in C++ with curllib with help from Petr
  • Test the 4D C++ interface
  • Working on DX AdSec flat issue
  • 4D meeting and demontration
This week
  • Contiue working on 4D
  • Change AO SW and test the a better way to determ jumping actutor with Guido
  • Improve the magnet detach prodcedure?
  • Follow SOUL commissioning?

YZ

  • LBC sky flat wiki program:
    • Finished a baseline version. Last Friday's and yesterday's results showed that it was successful.
    • Several refinements to be made.
    • Working on releasing an initial version for github.
    • Create a generic user account for editing the wiki page?
  • TCS:
    • Discussed with Petr, determined that several more changes to be made.
    • Working on fixing issues in the current pull request.
    • Submitting another pull request after closing the current one.
  • Start investigating on a generic infrastructure for instrument monitoring and alarm handling.

AC

  • AGW SSD planning
    • AGW1 removal July 18th. Estimated one week of work with engineering group for HK upgrade. Estimated one week of work with SWG for guide probe staleness.
  • AGW script to get current and following error telemetry
  • web-dms
  • test wda script
  • safety training
next:
  • continue wda, web-dms, AGW telemetry
  • OVMS+TT

SCPH

  • work on supportscripts (porting), packaging, etc; work on Steve A's scripts the same.
  • xephem for LBTO (Doug Miller's changes to xephem)
  • iLocater questions (Jonathan Crass on the mountain).
next:
  • calculate and purchase for R3 measurements taken two weeks ago.
  • kickstart John's Godzillux to obs1 base, kickstart display NUC to CentOS 7
  • finish up with supportscripts (is RB_Science.py working?)
  • monitor mountain -- upgrade Zabbix (probably), upgrade OS, add some watch points.

IS

Last week
  • LUCI
    • Reverted fix for #7642 (RTD aladin coordinate reading on X/Y has 0.5 px mismatch)
    • finished fixed #7473.
    • Disable automatic configuration writing to disk by LUCI processes, in particular GUIs if an environment variable: LUCI_DISABLE_CONFIG_WRITING is present.
    • Also fixed issue #6944. Compensation mirror locking up.
             E.g. some thread initializes mylock as:
              final Lock mylock = new ReentrantLock ();
              ...
              // tryLock() returns true if it gets the lock, false otherwise
              if (mylock.tryLock() && mytest == test) try {
                ...
              } catch() {
                ...
              } finally {
                mylock.unlock();
              }
      
              vs
      
              if (mytest == test) {
                if (mylock.tryLock()) {
                  try {
                    ...
                  } catch () {
                    ...
                  } finally {
                    mylock.unlock();
                  }
                }
              }
            
  • OT
    • Went through all open OT issues with Michelle and Andrew Cardwell and got a plan and a timeline to fix them.
This week
  • LUCI
    • Stow LUCI code and environment to the /usr/local
    • Look at what's needed to save SRR mode fits cube.
    • get all LUCI scripts in git
  • OT
    • Start working on OT issue list.
  • PIT
    • Make sure all scripts are in git
    • clean up Italian spreadsheet script generation.
Topic revision: r11 - 03 Jun 2019, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback